A point source emits 30.0 W of sound isotropically (in all directions). A small microphone with an area of 0.750 cm2 is located 200 m away from the source.

(a) Find the sound intensity at the location of the microphone in W/m2 .

(b) Find the power intercepted by the microphone.

(c) Find the displacement amplitude sm of the sound wave at the position of the microphone taking the frequency of the sound to be 50.0 Hz.
Answer #1

(a) I = P/A

I = 30 / 4pi*2002

I = 5.97e-5 W/m2

------------------------------------------------------------

(b) P = IA

P = 5.97e-5 * 7.5e-5

P = 4.48e-9 W

------------------------------------------------------------------

(c) For this part, I need to know velocity and density

the relation between displacement amplitude and intensity is

d = I / 22 * density * velocity * frequency2